OpenStudy (anonymous):

best anwser gets medal.. Highway construction crews often use the "grade" or steepness of a road in many of their calculations. A grade of 3% means that the road surface rises 3 feet for every horizontal run of 100 feet. A roadway that slopes up has a positive grade; likewise a roadway that slopes down has a negative grade. 1. If the vertical rise of a highway is 300 feet when the horizontal run is 5000 feet, what is the grade of the road? 2. If the grade of a road is 5%, what is the vertical rise of the road when the horizontal run is 4000feet

OpenStudy (anonymous):

3 On a certain stretch of road, the point 420 feet east of point A is 60 feet lower than point A. Point B, 600 feet east of A, is 84 feet lower than point A. a. Find the grade of the road surface between points A and B. b. Would a motorist driving along a stretch of road from point B to point A report the same grade for the road as a motorist driving from A to B would report? Explain your response.

OpenStudy (anonymous):

1. 3/100 = 3%. 300/5000 = ? 2. What is 5% of 4000? Or 0.05*4000 = ?

OpenStudy (anonymous):

what?